OpenStudy (anonymous):

Will Give medal! Evaluate. 3 x 1.5 - 2.5 ______________ 4 (8.62 - 6.37)

OpenStudy (idealist10):

\[\frac{ 4.5-2.5 }{ 4(2.25) }=\frac{ 2 }{ 9 }\]

OpenStudy (idealist10):

Answer: 2/9